How do you write integral denominators, and get the solution.
1/x+x/1.5=7/3

Answer by jorel1380(3719)

1/x+x/1.5=7/3
3/x+2x=7
3+2x2=7x
2x2-7x+3=0
(2x-1)(x-3)=0
x=1/2 or 3..